OpenStudy (anonymous):

What are the solutions to the equation x2 = 81?

OpenStudy (anonymous):

\[x^2=9^2\]

OpenStudy (anonymous):

ok cool! thanks

OpenStudy (anonymous):

its gonna be +9 and -9 cause when you find the sqrt there are two solutions a positive and a negative

OpenStudy (anonymous):

ok that is one of the answers so is only 9

OpenStudy (anonymous):

9*9=81 (-9)(-9)=81 as well

OpenStudy (anonymous):

soo is it -9 and 9

OpenStudy (maheshmeghwal9):

ya!

OpenStudy (anonymous):

yea but if there isnt an answer choice like that dont worry about it for now. ask your teacher though and look smart

OpenStudy (maheshmeghwal9):

\[Formula:-> x^2=a^2\]

OpenStudy (maheshmeghwal9):

Then, \[x= \sqrt{a^2}=\pm a\]